OpenStudy (anonymous):

I have NO idea! :( anyone? The arithmetic mean (A) of two numbers (a and b) is given by the formula A=a+b/2, and their geometric mean (G) is given by G=√ab. Their harmonic mean (H) is given by the formula G= √AH. Which formula correctly gives H in terms of a and b?

OpenStudy (anonymous):

H=2ab/a+b H=sqrt2ab/a+b H=a+b/2ab H=ab/a(a+b)

OpenStudy (anonymous):

Fan and medal who ever can help im desperate

zepdrix (zepdrix):

\[\Large\rm \color{royalblue}{A=\frac{a+b}{2}},\qquad\color{orangered}{G=\sqrt{ab}}\] And then,\[\Large\rm \color{orangered}{G}=\sqrt{\color{royalblue}{A}H}\]So we need to solve for H, the harmonic mean, ya? It's stuck under a square root, what should be our first step to get rid of the root?

OpenStudy (anonymous):

im back sorry.

OpenStudy (anonymous):

I've been trying to figure it out. Hope you come back soon.

zepdrix (zepdrix):

so, first step? :D

OpenStudy (anonymous):

I guess I would have to rewrite it H=G^2/A

zepdrix (zepdrix):

\[\Large\rm H=\frac{(\color{orangered}{G})^2}{\color{royalblue}{H}}\]Ok good so far.

OpenStudy (anonymous):

isolate lol

OpenStudy (anonymous):

im super slow and have total malfunctions when it come to numbers sorry

zepdrix (zepdrix):

\[\Large\rm H=\frac{(\color{orangered}{G})^2}{\color{royalblue}{H}}=\frac{\left(\color{orangered}{\sqrt{ab}}\right)^2}{\color{royalblue}{\frac{a+b}{2}}}\]Do you understand this step? I'm replacing G, with what it is, and replacing H, with what it is.

OpenStudy (anonymous):

oh yes :0 one question sqrtab^2 ...why did you square?

OpenStudy (anonymous):

oh nevermind i get it

zepdrix (zepdrix):

:3

zepdrix (zepdrix):

hopefully the colors help. I'm not doing anything too fancy. I'm just replacing the orange with the orange and the blue with the blue. you determined that the top is being squared.

OpenStudy (anonymous):

so now I have to solve

zepdrix (zepdrix):

so umm.. ya simplify. I guess we could do some `keep change flip` business, ya?

OpenStudy (anonymous):

give me a sec

zepdrix (zepdrix):

If you remember that little trick :) it tells us how to deal with fractions when we're dividing by them

OpenStudy (anonymous):

now im lost what trick?

OpenStudy (anonymous):

turn it upside down

zepdrix (zepdrix):

\[\Large\rm \frac{\left(\frac{a}{b}\right)}{\left(\frac{c}{d}\right)}=\frac{a}{b}\times\frac{d}{c}\]Yes good, flip the bottom fraction, turn the operation into multiplication.

zepdrix (zepdrix):

Works the same way if the top is not a fraction:\[\Large\rm \frac{a}{\left(\frac{c}{d}\right)}=a\times\frac{d}{c}\]Just need the bottom to be a fraction to apply this trick.

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

yeah im stuck on how to flip it being squared

OpenStudy (anonymous):

is this the hardest part lol

zepdrix (zepdrix):

Ok let's simplify the top before we do any flipping then :) What happens when you square a square root?

OpenStudy (anonymous):

multiply

zepdrix (zepdrix):

\[\Large\rm H=\frac{\sqrt{ab}^2}{\left(\frac{a+b}{2}\right)}\]

zepdrix (zepdrix):

Square root and square are inverse operations of one another, so they will `undo` each other.

zepdrix (zepdrix):

\[\Large\rm H=\frac{ab}{\left(\frac{a+b}{2}\right)}\]So this will happen to our top.

zepdrix (zepdrix):

Uh oh, what's wrong Molly? :o Confused?

OpenStudy (anonymous):

no you so much quicker then me

OpenStudy (anonymous):

you're

OpenStudy (anonymous):

oh yes I'm following you now

zepdrix (zepdrix):

Do you see how we're going to flip the (a+b)/2?

OpenStudy (anonymous):

2/ab

OpenStudy (anonymous):

no

OpenStudy (anonymous):

2/(a+b)

zepdrix (zepdrix):

\[\Large\rm H=\frac{ab}{\left(\frac{a+b}{2}\right)}\]The trick is called `keep` `change` `flip`. We `keep` the top the same. We `change` the operation from multiplication to division. We `flip` the bottom fraction. Ok good, our flip will look like that! So we get,\[\Large\rm H=ab\times\frac{2}{a+b}\]

OpenStudy (anonymous):

H=a+b/2ab

zepdrix (zepdrix):

Oo something got mixed up there :O

OpenStudy (anonymous):

see U are fast and I was wrong

zepdrix (zepdrix):

sorry i gotta slow down hehe

zepdrix (zepdrix):

your flip looked good though: 2/(a+b) is what it turned into, ya?

zepdrix (zepdrix):

Take a minute to soak all that in >.<

OpenStudy (anonymous):

wow! thank you so much for helping me . :)

zepdrix (zepdrix):

Do you understand the final step? We're so close! :O

OpenStudy (anonymous):

It will probably take some practice but I wrote it down.

OpenStudy (anonymous):

is my answer H=2ab/a+b

zepdrix (zepdrix):

Good good good. The ab will multiply into the numerator. It should be written like this though 2ab/(a+b) which is, I assume, what you meant :)

OpenStudy (anonymous):

yes woohoo :) this openstudy is fun and I hate math :)

zepdrix (zepdrix):

pshhh math is awesome c:

zepdrix (zepdrix):

Ohhh you're kinda new huh? Well then, \(\Large\bf \color{#CC0033}{\text{Welcome to OpenStudy! :)}}\)

OpenStudy (anonymous):

haha ok for you smart ones not me. takes me to long. I really appreciate it. Will you be on here this weekend? Can I ask for help again? I have to finish my review packet before monday exam. :(

zepdrix (zepdrix):

Ya, I'll be off and on at different times. Lots of other smart people on as well. You can sent like a "page" or notification to a specific person by typing @(their name) somewhere in the comments of your thread. Example @zepdrix If you look in the lobby, the people with 99 scores tend to be some smarty pants. So if I'm not around, you can try asking one of them for assistance as well :)

OpenStudy (anonymous):

ok thank you? :)

zepdrix (zepdrix):

This thread is going to get kinda slow and laggy though because of all the LaTeX I used. If you have another question ready, you should probably close this one and start a new one. np \c:/ hope the packet goes well!

OpenStudy (anonymous):

no not tonight but maybe tomorrow. Most of the stuff/questions I know but there are some things that I have never seen and not even sure what they are... like this question. I had never seen. Is that Algebra?

OpenStudy (anonymous):

I had a trig question too. Think my teacher hates me. lol

zepdrix (zepdrix):

Yah that's pretty much smack dab right in the middle of Algebra :) It's an algebraic expression that requires a bunch of simple steps (multiplication,division,etc) and some substitution.

OpenStudy (anonymous):

I had never heard of harmonic mean.lol simple :) thanks again good night

OpenStudy (chaosguardian14):

usein Tex commands looks like what u would type in a minecraft command block lol